OpenStudy (anonymous):

How do you calculate 6 to the power of -3 - it's part of my sons homework and I'm not sure how to teach him to do it

jimthompson5910 (jim_thompson5910):

6 to the power of 3 = 6*6*6 = 36*6 = 216

OpenStudy (vishweshshrimali5):

\[6^{-3} = \frac{1}{6}*\frac{1}{6}*\frac{1}{6}\]

OpenStudy (vishweshshrimali5):

\[= \frac{1}{216}\]

OpenStudy (anonymous):

Thank you!!!! This helps so much!!!!

jimthompson5910 (jim_thompson5910):

if the exponent is negative, then flip the fraction So 6^(-3) = 1/(6^3) = 1/(6*6*6) = 1/216

OpenStudy (anonymous):

so as a decimal how would it be written? 0.216?

jimthompson5910 (jim_thompson5910):

is the exponent positive or negative?

OpenStudy (anonymous):

for example 7 to the power of -2 is the answer 0.49?

jimthompson5910 (jim_thompson5910):

ok, the exponent is negative 7 to the power of -2 = 7^(-2) = 1/(7^2) = 1/49 = 0.0204081632653 I used a calculator to find that last part

jimthompson5910 (jim_thompson5910):

6 to the power -3 = 6^(-3) = 1/(6^3) = 1/216 = 0.00462962962962

OpenStudy (anonymous):

The question on his home work is simplify each power - and it is 7 to the power of -2 - so I guess the answer would be 1/7*1/7 right?

jimthompson5910 (jim_thompson5910):

1/7*1/7 = 1/(7*7) = 1/49 So the answer is 1/49

jimthompson5910 (jim_thompson5910):

I'd leave it at that and avoid the decimal part of the answer

OpenStudy (anonymous):

yeah I think you're right about the decimal part

OpenStudy (anonymous):

thank you